Calculate for; Mary got 58% in an examination, scoring 348 marks, her sister scored 396 marks in the same examination. Calculate her sister's percentage score
Answer by TimothyLamb(4379)

58/348 = x/396
x = (58/348)*396
x = 66%
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php